The measure of an angle is 20° more than he measure of its complement. What is the smaller angle?
Answer by oberobic(2304)

x = angle
90-x = its complement
x-20 = 90-x
2x = 110
x = 55
90-x = 35
Answer: The smaller angle is 35 degrees.